iF x varies inversely as y^2 and x=3 when y=24, find x when y=6
Answer by fractalier(6550)

If x varies inversely as y^2, we write that relationship as
x = k / y^2
Now plug in the original data to find the k...
3 = k / 24^2
k = 3 * 576 = 1728
Now rewrite the relationship, using this k and the new data to find x...
x = 1728 / y^2
x = 1728 / 6^2
x = 1728 / 36
x = 48